According to statistics reported on CNBC, a surprising number of motor vehicles are not covered by insurance. Sample results, consistent with the CNBC report, showed 30 of 203 vehicles were not covered by insurance a. What is the point estimate of the proportion of vehicles not covered by insurance (to 4 decimals)? b. Develop a 90% confidence interval for the population proportion (to 4 decimals)
Answer #1

Answer

(A) Point estimate of the proportion of vehicles not covered by insurance= (vehicles not covered by insurance)/(total vehicles)

setting the given values from the question

we get

Point estimate = 30/203 = 0.1478 (rounded to 4 decimals)

(B) Confidence interval formula for the population proportion is given as

CI = hat{p} pm z*sqrt{(hat{p}*(1-hat{p}))/n}

where p = 0.1477, n = 203

Using z distribution table, the critical z value for 90% confidence interval is 1.64

setting the given values, we get

C1-0.1477 ± 1.64 * V/(0.1477 * (1-0.1477)/203

on solving, we get CI 0.1477士1.64 * 0.0249 = (0.1069. 0. 1885)

So, the required 90% confidence interval is (0.1069,0.1885)